OpenStudy (anonymous):

Solve each system using substitution. Check your answer. y=8-x 7=2-y

Mehek (mehek14):

7 = 2 - 8 -x 7 = -6 - x 13 = - x -13 = x

OpenStudy (anonymous):

thank you